OpenStudy (gabylovesyou):

When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m?

OpenStudy (gabylovesyou):

is it 3 ?

jimthompson5910 (jim_thompson5910):

what do you get when you solve for y

OpenStudy (gabylovesyou):

y = 3(x - 8) / 4 ??

jimthompson5910 (jim_thompson5910):

can you write that in the form y = mx+b

jimthompson5910 (jim_thompson5910):

and what you got is a bit off, but you're onto the right path to get the correct slope

OpenStudy (gabylovesyou):

hmm...

OpenStudy (gabylovesyou):

idk

jimthompson5910 (jim_thompson5910):

Here's a similar example \[\large 2x - 7y = 35\] \[\large -7y = -2x+35\] \[\large y = \frac{-2x+35}{-7}\] \[\large y = \frac{-2x}{-7}+\frac{35}{-7}\] \[\large y = \frac{2}{7}x-5\] The last equation is in y = mx+b form where m = 2/7 b = -5 So the slope of this example is 2/7 This is NOT your answer as it's just an example

OpenStudy (gabylovesyou):

3x - 4y = 24 -4y = -3x + 24 y = -3x + 24 / -4 y = -3x / -4 + 24 / -4 y = 3/4x - 6

jimthompson5910 (jim_thompson5910):

perfect, you nailed it

jimthompson5910 (jim_thompson5910):

so the slope is m = ???

OpenStudy (gabylovesyou):

3/4

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (gabylovesyou):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome